# Alternant Hydrocarbon Diradicals as Optically Addressable Molecular Qubits

### 14 Mar 2024 v1

@@ Abstract
High-spin molecules allow for bottom-up qubit design and are promising platforms for magnetic sensing and quantum information science. Optical addressability of molecular electron spins has also been proposed in first-row transition metal complexes via optically-detected magnetic resonance (ODMR) mechanisms analogous to the diamond-NV colour centre. However, significantly less progress has been made on the front of metal-free molecules, which can deliver lower costs and milder environmental impacts. At present, most luminescent open-shell organic molecules are $\pi$-diradicals, but such systems often suffer from poor ground-state open-shell characters necessary to realise a stable molecular qubit. In this work, we propose the use of alternancy symmetry to selectively minimise radical-radical interactions in the ground state, generating $\pi$-systems with high diradical characters. We call them m-dimers, referencing the need to covalently link two benzylic radicals at their meta carbon atoms for the desired symmetry. Through a detailed electronic structure analysis, we find that the excited states of alternant hydrocarbon m-diradicals contain important symmetries that can be used to construct ODMR mechanisms. The molecular parameters are set in the context of a tris(2,4,6-trichlorophenyl)methyl (TTM) radical dimer covalently tethered at the meta position, demonstrating the feasibility of realising a molecular colour centre with alternant $\pi$-diradicals.
@@ I Introduction
There has been sustained interest in designing optically addressable electron spins that can be both initialised (polarised) and read out by optical means [1]. Known as optically-detected magnetic resonance (ODMR), this technique offers precise control of a single qubit and has found numerous applications in quantum sensing [2] and quantum information science [3]. The platform, referred to as a colour centre, is often a solid-state spin defect such as a nitrogen-vacancy (NV) centre in diamond [4, 5, 6, 7, 8]. However, because these defects are introduced post-synthesis and often without control over their locations, defect-based colour centres suffer from poor scalability and tunability.
These problems can be tackled with molecular spin systems because they can be synthesised from bottom-up and extended into macromolecular systems [9, 10, 11, 12, 13]. Metal complexes have seen the most advancements, particularly with transition metals [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. By contrast, little progress has been made with fully organic molecules [26, 27, 28, 29, 30] which, being metal-free, will be more cost-efficient and sustainable than their metal-based counterparts. Following the pioneering work of Gorgon et al. in synthesising an optically-addressable organic colour centre [27], we present the first theoretical analysis on the spin-optical interface of such systems.
In diamond-NV centres, ODMR is achieved via the following procedure [31]: Upon photoexcitation of the triplet ground state to the excited state (also a triplet), intersystem crossing (ISC) to a singlet excited state occurs with a change in magnetic quantum number $M_{S}$. This increases the relative population of the excited triplet $M_{S}=0$ level, the photoluminescence (PL) of which transfers the spin polarisation to the ground state (optical initialisation). The PL intensity also indicates the ground state polarisation (optical readout). As for the singlet excited state, its population decays non-radiatively to the triplet $M_{S}=0$ ground state, its irreversibility being the key towards overall purification of the ground state spins [Fig. Ia]. There are thus three elements to engineer in a metal-free molecular system: (1) A high-spin ground state, (2) spin-selective interactions among the excited states, and (3) irreversible relaxation to the ground state.
@@
Element (1) suggests the use of an organic diradical. Thus far, most luminescent diradicals have involved some degree of $\pi$-conjugation [32, 33, 34, 35, 36, 37, 38, 39], which has proven useful in the chemical assembly of spin arrays [33]. However, large spatial separation between the two paramagnetic centres is often necessary to minimise radical-radical interactions and achieve a high-spin ground state [Fig. Ib] [40, 41, 42, 43]. In most cases, this separation extends to the excited states as well, which is disadvantageous for element (2). Indeed, in Gorgon et al.’s seminal work, interradical interaction was achieved via a charge transfer to the spacer’s triplet state, generating a complicated manifold of 12 coupled electronic levels [27]. To avoid this issue, future designs may reduce (improve) the stability of the molecule’s closed-shell (open-shell) Kekulé structures, with a crude measure being the number of Clar sextets minus the number of unpaired electrons [44]. In some extreme cases, $\pi$-diradicals may have no closed-shell Kekulé structures at all; such molecules are termed “non-Kekulé” [40, 43] and an example has been shown in Fig. Ic. However, this analysis is purely qualitative and there exist non-Kekulé molecules with appreciable closed-shell characters [42]. (Interestingly, alkaline earth metal complexes [45], commonly used as optical cycling centres [46, 47], as well as nitrenes [48] have also been investigated for luminescent diradical properties and are useful candidates of molecular colour centres as well.)
@@
Therefore, through a qualitative analysis, we find potential in realising an NV centre analogue by attaching two AHRs covalently at selected atoms and with significant torsion. We call them AHR m-dimers, referencing the fact that most luminescent $\pi$-radicals have benzylic structures and thus should be dimerised via atoms at the meta position. In this work, we make further progress by analysing the electronic structure of AHR m-dimers and the symmetries of their excited wavefunctions. By adopting a mean-field approach to the Pariser-Parr-Pople (PPP) model [55, 56, 57] (also known as the extended Hubbard model), we find that not all excitations are emissive and only few within the subset can undergo triplet-singlet ISC. This makes the excited state kinetics highly selective and useful for generating robust ODMR mechanisms. These observations are set in the context of two existing AHRs – a methylated benzylic radical and the tris(2,4,6-trichlorophenyl)methyl (TTM) radical – using parameters estimated from density functional theory (DFT), further demonstrating the possibility of designing a molecular colour center with AHR m-dimers. For the methylated benzylic radical, our DFT-parametrised excited states also matched the results from a multi-configurational level of theory.
The paper is organised as such: In Sec. IIA, the electronic structure of AHR m-dimers is analysed using the PPP framework. This section is technical and the reader may safely skip it at the first pass. In Sec. IIB, ODMR mechanisms are proposed based on results from the PPP model. In Sec. IIC, the PPP predictions are verified by ab initio calculations. Finally, in Sec. IID, a mapping between a polarised ground-state triplet and an operational qubit is proposed.

### 18 Mar 2024 v2

@@ Abstract
High-spin molecules allow for bottom-up qubit design and are promising platforms for magnetic sensing and quantum information science. Optical addressability of molecular electron spins has also been proposed in first-row transition metal complexes via optically-detected magnetic resonance (ODMR) mechanisms analogous to the diamond-NV colour centre. However, significantly less progress has been made on the front of metal-free molecules, which can deliver lower costs and milder environmental impacts. At present, most luminescent open-shell organic molecules are $\pi$-diradicals, but such systems often suffer from poor ground-state open-shell characters necessary to realise a stable ground-state molecular qubit. In this work, we use alternancy symmetry to selectively minimise radical-radical interactions in the ground state, generating $\pi$-systems with high diradical characters. We call them m-dimers, referencing the need to covalently link two benzylic radicals at their meta carbon atoms for the desired symmetry. Through a detailed electronic structure analysis, we find that the excited states of alternant hydrocarbon m-diradicals contain important symmetries that can be used to construct ODMR mechanisms leading to ground-state spin polarisation. The molecular parameters are set in the context of a tris(2,4,6-trichlorophenyl)methyl (TTM) radical dimer covalently tethered at the meta position, demonstrating the feasibility of alternant m-diradicals as molecular colour centres.
alternant hydrocarbon diradicals, molecular qubits, metal-free, ground-state spin polarisation, molecular colour centres
@@ I Introduction
There has been sustained interest in designing optically addressable electron spins that can be both initialised (polarised) and read out by optical means [1]. Known as optically-detected magnetic resonance (ODMR), this technique offers precise control of a single qubit and has found numerous applications in quantum sensing [2] and quantum information science [3]. The platform, referred to as a colour centre, is often a solid-state spin defect such as a nitrogen-vacancy (NV) centre in diamond [4, 5, 6, 7, 8, 9]. However, because these defects are introduced post-synthesis and often without control over their locations, defect-based colour centres suffer from poor scalability and tunability.
These problems can be tackled with molecular spin systems because they can be synthesised from bottom-up and extended into macromolecular systems [10, 11, 12, 13, 14, 15, 16]. Metal complexes have seen the most advancements, particularly with transition metals [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. By contrast, little progress has been made with fully organic molecules [29, 30, 31, 32, 33, 34] which, being metal-free, will be more cost-efficient and sustainable than their metal-based counterparts. Pioneering efforts were made by Gorgon et al. using excited organic radicals [31]; however, a drawback of microwave spin manipulation in an excited electronic state is the limited lifetime of the latter. As for ground-state organic systems, the same (and only) work [31] also showed light-induced selective preparation of a diradical in the triplet ground state over its degenerate singlet counterpart. In this case, it is unclear if the triplet magnetic sublevels were differentially populated, which is key towards coherent spin control. To this end, we present the first theoretical analysis for a general class of organic spin-optical interfaces, focusing on attaining spin polarisation in the ground-state magnetic sublevels.
In diamond-NV centres, ODMR is achieved via the following procedure [35]: Upon photoexcitation of the triplet ground state to the excited state (also a triplet), intersystem crossing (ISC) to a singlet excited state occurs with a change in magnetic quantum number $M_{S}$. This increases the relative population of the excited triplet $M_{S}=0$ level, the photoluminescence (PL) of which transfers the spin polarisation to the ground state (optical initialisation). The PL intensity also indicates the ground-state polarisation (optical readout). As for the singlet excited state, its population decays non-radiatively to the triplet $M_{S}=0$ ground state, its irreversibility being the key towards overall purification of the ground state spins [Fig. Ia]. There are thus three elements to engineer in a metal-free molecular system: (1) a high-spin ground state, (2) spin-selective interactions among the excited states, and (3) irreversible relaxation to the ground state.
@@
Element (1) suggests the use of an organic diradical. Thus far, most luminescent diradicals have involved some degree of $\pi$-conjugation [36, 37, 38, 39, 40, 41, 42, 43], which has proven useful in the chemical assembly of spin arrays [37]. However, large spatial separation between the two paramagnetic centres is often necessary to minimise radical-radical interactions and achieve a high-spin ground state [Fig. Ib] [44, 45, 46, 47]. In most cases, this separation extends to the excited states as well, which is disadvantageous for element (2). Indeed, in Gorgon et al.’s seminal work, interradical interaction was achieved via a charge transfer to the spacer’s triplet state, generating a complicated manifold of 12 coupled electronic levels [31]. One solution would be to reduce (improve) the stability of the molecule’s closed-shell (open-shell) Kekulé structures, with a crude measure being the number of Clar sextets minus the number of unpaired electrons [48]. In the most extreme case, the $\pi$-diradical may have no closed-shell Kekulé structures at all; such molecules are termed “non-Kekulé” [44, 47] and an example has been shown in Fig. Ic. However, this analysis is purely qualitative and there exist non-Kekulé molecules with appreciable closed-shell characters [46]. (Interestingly, alkaline earth metal complexes [49], commonly used as optical cycling centres [50, 51], as well as nitrenes [52] have also been investigated for luminescent diradical properties and are useful candidates of molecular colour centres as well.)
@@
Therefore, through a qualitative analysis, we find potential in realising an NV centre analogue by attaching two AHRs covalently at selected atoms and with significant torsion. We call them AHR m-dimers, referencing the fact that most luminescent $\pi$-radicals have benzylic structures and thus should be dimerised via atoms at the meta position. In this work, we make further progress by analysing the electronic structure of AHR m-dimers and the symmetries of their excited wavefunctions. By adopting a mean-field approach to the Pariser-Parr-Pople (PPP) model [59, 60, 61] (also known as the extended Hubbard model), we find that not all excitations are absorptive and only few within the subset can undergo triplet-singlet ISC. This makes the excited state kinetics highly selective and useful for generating robust ODMR mechanisms leading towards ground-state spin polarisation. While general to any AHR m-dimer, our observations are also set in the context of two existing AHRs – a methylated benzylic radical and the tris(2,4,6-trichlorophenyl)methyl (TTM) radical – using parameters estimated from density functional theory (DFT). For the methylated benzylic radical, our DFT-parametrised excited states also matched the results from a multi-configurational level of theory. These ab initio calculations further demonstrate the possibility of designing a molecular colour center with AHR m-dimers.
The paper is organised as such: In Sec. IIA, the electronic structure of AHR m-dimers is analysed using the PPP framework. This section is technical and the reader may safely skip it at the first pass. In Sec. IIB, ODMR mechanisms resulting in ground-state spin polarisation are proposed based on results from the PPP model. In Sec. IIC, the PPP predictions are verified by ab initio calculations. Finally, in Sec. IID, a mapping between a polarised ground-state triplet and an operational qubit is proposed.